OpenStudy (anonymous):

2. solve. 4x + 6 < -6

OpenStudy (anonymous):

4x + 6 < -6 -6 - 6 = 12 4x = -12 -12 / 4 = ? can you do that last part?

OpenStudy (anonymous):

yeah, thanks!

OpenStudy (anonymous):

no problem ! :)